# Velocitune: A Velocity-based Dynamic Domain Reweighting Method for Continual Pre-training

### v1

@@ 3.2 Reasoning training results
Velocitune learns math and keeps coding ability.
In Table 1, we list the the nine benchmarks tested for comparing the methods on CodeLlama-7B with Reasoning. Velocitune leads the CPT baseline by 1.9% on seven math benchmarks and 3.8% on code average while baseline’s coding ability dropped by 1.3% from before the continued pre-train. It underscores the reason behind Velocitune’s effectiveness, is balancing the learning velocity in each domain, while in the baseline due to the static domain mixture, the model might learn some domains very well while not saturate for other domains. Velocitune by align the learning velocity across domains, resulting in a more balanced learning progress, thus developing a model better in downstream tasks.
Figure 3: Domain weights dynamic of Velocitune and DBL in training Llama-3 8B on SystemStack.
Table 2: Results of Velocitune, DBL, and basline on Llama3 and Mistral on CmdGen-NVIDIA and CmdGen-AMD benchmarks.
@@
Velocitune dynamically regularises learning velocity.
As shown in Figure 2, the overall weight adjustments can be divided into three distinct stages. During the initial phase of training, the weights for newly introduced domains—OpenWebMath, Arxiv, and AlgebraicStack—rise, while those for the replay domains, GenCode and GenLanguage, decline sharply. This occurs because the LMs are typically less saturated in the new domains, leading to slower learning compared to the replay domains. As a result, Velocitune increase the weights of the underperforming domains to balance their learning. In the second stage, the weights for the replay domains begin to increase while those of the new domains decrease. This shift happens after the models have made substantial progress in the new domains, prompting Velocitune to reallocate focus toward the replay domains. In the final stage, the replay domains’ weights decrease again, while the downslopes for AlgebraicStack and OpenWebMath become more gradual, even showing a slight peak before continuing to decline. This suggests that the model has learned these domains well. Meanwhile, Velocitune increases the weight of Arxiv, where the learning velocity remains slower compared to the others. By adjusting dynamically, Velocitune aligns the learning velocity across domains, ensuring a more balanced learning outcome.
@@ 3.3 SystemStack training results
Table 3: Performance comparison of CodeLlama-7B trained using reweighted domain ratios, original domain ratios,Velocitune, and Velocitune without target loss on math and code reasoning benchmarks.
@@
Table 4: Performance comparison of Llama3 models trained reweighted domain ratios, original domain ratios, and Velocitune on the CmdGen benchmarks
@@
On SystemStack, Velocitune improves downstream task performance over DBL and CPT on Llama3-8B and Mistral-7B. Table 2 shows that Velocitune improves the performance in CMD-NVIDIA by 5.3% and 4.4% compared to baseline on Llama3-8B and Mistral-7B, and 3.1% and 2.6% in CMDGen-AMD. It is interesting to find that on Llama3-DBL underperforms baseline in both benchmarks which could attribute to an imbalanced learning brought by updating domain weights based on the distances to the target loss.
Velocitune accelerates weight stablisation
Velocitune accelerates weight stablisation for system domain. We plot the weights moving trajectory on SystemStack comparing Velocitune, DBL, and the baseline in Figure 3. In parallel with Xia et al. (2023), we also observe that the weights stabilize after a few thousand steps. As re-scaling accelerates convergence in machine learning (Juszczak et al., 2002), we also observe that Velocitune on SystemStack reaches stablisation at least 1.5x faster than DBL as DBL does not stablise at the end of training. Observing from the trend, the curves of DBL seem to be a prolonged version of Velocitune. We hypothesis that if the DBL training continues for more steps, it will reach the same final weights similar to Velocitune. For training on Reasoning, we do not observe a stablisation but only the curve becoming towards flat at the end of training.
3.4 Data ordering contribute along with reweight domain weights
To isolate the key contributors behind the effectiveness of Velocitune, we further conduct an ablation study. Specifically, we collect the weights during Velocitune for every evaluation interval, then average over the weights to have the overall sampling ratio of each domain across the training. A comparison of the originial weight and the reweighted weights are shown in Table 9 and 10. Provided with the reweighted data ratio, we train the model on the dataset constantly on the mixture for the same training steps to examine its effect. We conduct the experiment for CodeLlama-7B on Reasoning and Llama3 on SystemStack. The results are shown in Table 3 and Table 4.
Table 5: Performance comparison of CodeLlama-7B with or without target loss.
@@ 3.4 Data ordering contribute along with reweight domain weights
On the CmdGen benchmarks, the Llama-3 trained with reweighted ratio data outperform Llama3-CPT on both benchmarks, but inferior to Velocitune on NVIDIA and tie on AMD. On math and code evaluation, the CodeLlama trained with reweighted ratio achieves the same average score with Velocitune on the math benchmarks, but still fall behinds Velocitune by 2.7% for coding benchmarks. In general, reweighted data ratios improve downstream task performance compared to the original mixture. However, models still underperform relative to Velocitune, with the sole exception being the Math average, where performance is 0.7% higher than Velocitune. This finding is particularly intriguing, as data mixture is typically assumed to be the primary factor influencing downstream performance. However, this result emphasizes the critical role of dynamic data mixing rather than relying on a static mixture. Previous studies (Hu et al., 2024; AI@Meta, 2024) have highlighted the effects of presenting different data at various stages of model training. Our comparison further supports the potential of dynamic data weighting during pre-training. We encourage further research into the impact of data ordering and its role in optimizing model performance.
@@ 3.5 Target loss is critical to Velocitune
To highlight the significance of the target loss, we also explore an alternative method of updating the weights by re-scaling the evaluation losses during training. Specifically, we divide the evaluation loss by the initial loss for each domain, as given by: $\delta^{i}=\frac{\ell^{i}_{T}}{\ell^{i}_{\text{init}}}$. The domain weight dynamic of this is shown in Figure 4 and the corresponding results are presented in Table 5. We see that without the constraints of target loss, scaling solely on the initial loss constantly drives the weight of GenLangauge to nearly 1 and surpress others to 0. Consequently, the model exhibits a performance decrease of 4.1% in the math benchmarks and 0.5% in the code benchmarks compared to the CPT baseline. We hypothesize that this update function implicitly assumes a target loss of zero for each domain, leading to unrealistic expectations for the model to minimize the loss to zero across all domains. Such impractical targets disrupt the training process, ultimately degrading the model’s overall performance.
@@ 4 Related Work
GLaM (Du et al., 2022) manually assigns weights to each domain based on the amount of data and the downstream performance of a small model trained on the domain data. Lemma (Azerbayev et al., 2023), an effort to continue pretraining LMs with math-related data from three domains, determines domain weights by iterating through different combinations and selecting those with the lowest perplexity on an evaluation dataset. Doremi (Xie et al., 2024) employs a three-stage process to dynamically adjust domain weights: (1) Train a reference model, (2) Train a proxy model while adjusting the proportion of data based on the proxy model’s loss, and (3) Train the final model using the aggregated data weights, demonstrating superior performance compared to models trained with the original weights. Xia et al. (2023) shows that instead of training a reference model, the target loss value can be estimated using scaling laws (Hoffmann et al., 2022), achieving better performance on smaller LMs with the predicted losses.
@@ 5 Conclusion
In this work, we introduced Velocitune, a novel approach for dynamically adjusting domain weights during the training of large language models. Our method addresses the limitations of existing techniques that rely on static or purely loss-based adjustments, which can lead to imbalanced learning across different domains. By scaling to align learning velocity across domains, Velocitune ensures a more balanced learning process, thereby enhancing the model’s ability to generalize across diverse tasks. Our experiments on CodeLlama 7B and Llama3 demonstrate the effectiveness of Velocitune in improving performance on a variety of tasks, including math reasoning, coding, and system diagnosis. Furthermore, our ablation study reveals that the benefits of Velocitune are not solely due to the adjustment of domain weights, but also stem from the synergistic effect of the training dynamics, including the order in which data from different domains is processed. This finding underscores the importance of considering both the weights and the order of learning in the pre-training of large language models.
@@ 6 Limitations
The limitations of our study are twofold. Firstly, Velocitune only applies to continual pre-training, and is yet to be examined in pre-training from scratch. Secondly, despite the multi-domain nature of prevelant supervised fine-tuning(SFT) datasets, the possibility of applying Velocitune in SFT has not been explored.
Appendix A Examples of CmdBen
@@
Appendix B Scaling law prediction error
@@ Appendix D Training Details
We ran the experiments on 64 Nvidia H100s.
@@ Appendix E Reweighted Data Ratio after Velocitune
The reweighted data ratio is shown in Table 3 and Table 9.
Table 9: Original and Velocitune reweighted ratio on SystemStack for Llama3
Domain Blogs Stackoverflow Arxiv
Original 0.187 0.503 0.310
@@
Table 10: Original and Velocitune reweighted ratio on Reasoning for CodeLlama
Domain AlgebraicStack Arxiv OpenWebMath GenCode GenLanguage
Original 0.189 0.500 0.259 0.029 0.020
Reweighted 0.185 0.463 0.225 0.086 0.040
Appendix F Domain weight dynamic in Velocitune without target loss

### v2

@@ 3.2 Reasoning training results
Velocitune learns math and maintains coding ability.
In Table 1, we list the benchmarks tested to compare the methods in CodeLlama-7B with Reasoning. Velocitune leads the baseline by 1.6% on average across eight math benchmarks and 3.8% on code average while the baseline’s coding performance dropped by 1.3% from before the continued pre-training. It underscores the reason behind Velocitune’s effectiveness, which is balancing the learning velocity in each domain, while in the baseline due to the static domain mixture, the model might learn some domains well while not saturating for other domains. Velocitune aligns the learning velocity across domains, resulting in more balanced learning progress, thus developing models of better downstream performance.
Figure 3: Domain weights dynamic of Velocitune and DBL in training Llama-3 8B on SystemStack. The vertical axis represents domain weights, while the horizontal axis denotes training steps.
Table 2: Results of Velocitune, DBL, and Basline on Llama3 and Mistral on CmdGen-NVIDIA and CmdGen-AMD benchmarks. Best results for each metric are highlighted in bold.
@@
Velocitune dynamically regularises the learning velocity.
As shown in Figure 2, the overall weight adjustments can be divided into three distinct stages. During the initial phase of training, the weights for newly introduced domains—OpenWebMath, Arxiv, and AlgebraicStack—rise, while those for the replay domains, GenCode and GenLanguage, decline sharply. This occurs because LMs are typically less saturated in the new domains, leading to slower learning compared to replay domains. As a result, Velocitune increases the weights of the underperforming domains to balance their learning. In the second stage, the weights for the replay domains begin to increase, while those of the new domains decrease. This shift occurs after the models have made substantial progress in the new domains, prompting Velocitune to reallocate the focus to the replay domains. In the final stage, the weights of the replay domains decrease again, while the downslopes for AlgebraicStack and OpenWebMath become more gradual, even showing a slight peak before continuing to decline. This suggests that the model has learnt these domains well. Meanwhile, Velocitune increases the weight of Arxiv. Adjusting weights dynamically, Velocitune aligns the learning velocity across domains, ensuring a balanced learning outcome.
@@ 3.3 SystemStack training results
Table 3: Performance comparison of CodeLlama-7B trained using Reasoning data with default domain ratios, reweighted domain ratios, and Velocitune, evaluated on math and code reasoning benchmarks.
@@
Table 4: Performance comparison of Llama3 models trained using SystemStack with original domain ratios, reweighted domain ratios, and Velocitune, evaluated on the CmdGen benchmarks
@@
On SystemStack, Velocitune improves downstream task performance over DBL and Baseline on Llama3-8B and Mistral-7B. Table 2 shows that Velocitune improves the Accuracy in CMD-NVIDIA by 4.9% and 4.4% compared to baseline on Llama3-8B and Mistral-7B, and 3.1% and 2.6% and in CMDGen-AMD. We also found that Llama3-DBL underperforms the baseline on both benchmarks, which could be attributed to an imbalanced learning brought by updating domain weights based on the distances to the target loss.
Velocitune accelerates weight stabilisation
We plot the movement trajectory of the weights in SystemStack comparing Velocitune, DBL, and the baseline in Figure 3. In parallel with Xia et al. (2023), we also observe that the weights stabilise after a few thousand steps. As re-scaling accelerates convergence in machine learning (Juszczak et al., 2002), we also observe that Velocitune on SystemStack reaches stabilisation at least 1.5x faster than DBL as DBL does not stabilise at the end of training. We hypothesise that if the DBL training continues for more steps, it will reach the same final weights as Velocitune. For training on Reasoning which is shown in Figure 2, we do not observe a complete stabilisation, but only the curve becoming flat at the end of training.
3.4 Data ordering contributes along with reweighted domain weights
To isolate the key contributors behind the effectiveness of Velocitune, we performed an ablation study to examine the effect of the reweighted data proportion. Specifically, we collect the weights during Velocitune for each evaluation interval and then average the weights to get the overall sampling ratio of each domain during training. A comparison of the original ratio and the reweighted ratio is shown in Tables 9 and 10. Using the reweighted data ratio, we train the model on the dataset using the static mixture for the same training steps. We conducted the experiment for CodeLlama-7B on the Reasoning dataset and Llama3 on SystemStack. The results are shown in Table 3 and Table 4.
Table 5: Performance comparison of CodeLlama-7B trained using Reasoning data with or without target loss and Baseline, evaluated on math and code benchmarks.
@@ 3.4 Data ordering contributes along with reweighted domain weights
On CmdGen benchmarks, Llama-3 trained with reweighted ratio data(denoted as Llama-3-Reweighted) outperforms Llama3-Baseline on both benchmarks, but is inferior to Velocitune on NVIDIA and on par with Velocitune on AMD. In the evaluation of math and code reasoning, CodeLlama trained with the reweighted ratio(denoted as CodeLlama-Reweighted) achieves a slightly higher accuracy than Velocitune on the math benchmarks, but still falls behind Velocitune by 2.7% for the coding benchmark. In general, reweighted data ratios improve downstream task performance compared to the original mixture. However, the models still underperform relative to Velocitune, with the sole exception being the Math average, where the performance is 0.7% higher than Velocitune. This finding sheds new light on the common assumption that data mixture is the primary factor driving downstream performance. Previous studies (Hu et al., 2024; AI@Meta, 2024) have highlighted the effects of presenting different data at various stages of model training. Our comparison further supports the potential of dynamic data weighting during pre-training. We encourage further research into the impact of data ordering and its role in optimising model performance.
@@ 3.5 Target loss is critical to Velocitune
To highlight the significance of the target loss estimate in learning velocity, we also explore an alternative method that excludes the target loss in learning velocity when updating the weights. In this setting, Equation 3 is reduced to dividing the evaluation loss by the initial loss for each domain, as given by: $\delta^{i}=\frac{\ell_{t}^{i}(\theta,S^{\prime}_{\text{eval},i})}{\ell^{i}_{\text{init}}}$. Using this formulation, we train CodeLlama on the Reasoning dataset. The evaluation results are presented in Table 5 and the dynamic of the domain weight is shown in Figure 4. We see that without the constraints of target loss, the weight of GenLangauge is driven to nearly 1 and others are suppressed to nearly 0. Consequently, the model exhibits a performance decrease of 4.1% in the math benchmarks and 0.5% in the code benchmarks compared to the baseline. We hypothesise that this update function implicitly assumes a target loss of zero for each domain, leading to unrealistic expectations for the model to minimise the loss to zero across all domains. Such impractical targets disrupt the training process, ultimately degrading overall performance.
@@ 4 Related Work
GLaM (Du et al., 2022) manually assigns weights to each domain based on the amount of data and the downstream performance of a small model trained on the domain data. Lemma (Azerbayev et al., 2023), an effort to continue pre-training LMs with math-related data from three domains, determines domain weights by iterating through different combinations and selecting those with the lowest perplexity on an evaluation dataset. Doremi (Xie et al., 2024) employs a three-stage process to dynamically adjust domain weights: (1) Train a reference model, (2) Train a proxy model while adjusting the proportion of data based on the proxy model’s loss, and (3) Train the final model using the aggregated data weights, demonstrating superior performance compared to models trained with the original weights. Xia et al. (2023) shows that instead of training a reference model, the target loss value can be estimated using scaling laws (Hoffmann et al., 2022), achieving better performance on smaller LMs with the predicted losses.
@@ 5 Conclusion
In this work, we introduced Velocitune, a novel approach to dynamically adjust domain weights during the training of large language models. Our method addresses the gap of existing works in domain-adapative continual pre-training. By aligning learning velocity across domains, Velocitune ensures a more balanced learning process, thereby enhancing the model’s ability to generalise across diverse downstream tasks. Our experiments on CodeLlama-7B, Llama3-8B, and Mistral-7B demonstrate the effectiveness of Velocitune in improving performance in a variety of tasks, including math reasoning, coding, and system command generation. Furthermore, our ablation study reveals that the benefits of Velocitune are not solely due to the adjustment of the domain weights, but also stem from the synergistic effect of the training dynamics, including the order in which data from different domains are processed. This finding underscores the importance of considering both the weights and the data order of in continual pre-training.
@@ 6 Limitations
The limitations of our study are threefold. First, Velocitune is currently designed for continual pre-training and has not yet been evaluated in pre-training from scratch. Second, while supervised fine-tuning (SFT) datasets often span multiple domains, the applicability of Velocitune in the SFT stage remains unexplored. Third, Velocitune incurs higher costs than the baseline due to its evaluation process during training. However, since pre-training primarily prioritises downstream performance, we defer a detailed analysis of computational efficiency and potential optimisations of Velocitune to future work.
7 Ethics Statement
This research is based on open-source models and is intended solely for research purposes. We acknowledge potential risks. While Velocitune improves model efficiency and performance, improved reasoning and code generation capabilities could be misused in misinformation generation, automated decision making, or bias reinforcement. Our work focusses on specific domains (maths, code, and system commands), which may not generalise well to under-represented languages or topics. Our research does not involve sensitive or private data.
Appendix A Examples of CmdGen
We show in the following two question and answer pairs from the CMDGen benchmarks.
@@ Appendix A Examples of CmdGen
Appendix B Scaling law prediction details
We continue training until the difference between the predicted target losses using the previous checkpoint sets and the current checkpoint set becomes smaller than a specified threshold (denoted $\sigma$ ). This criterion helps to select an appropriate value for k during the proxy model training. Detailed scaling law prediction errors are given in Table 6.
@@ Appendix D Training Details
We run the experiments on 64 Nvidia H100 GPUs. Our distributed training is based on accelerate [^4] https://huggingface.co/docs/accelerate/en/index and FDSP [^5] https://pytorch.org/blog/introducing-pytorch-fully-sharded-data-parallel-api/.
@@ Appendix E Reweighted Data Ratio after Velocitune
The reweighted data ratio is shown in Table 9 and Table 10
Table 9: The domain weights of Baseline and Velocitune’s reweighted domain weights on SystemStack for Llama3
Blogs Stackoverflow Arxiv
Baseline 0.187 0.503 0.310
@@
Table 10: The domain weights of Baseline and Velocitune’s reweighted domain weights on Reasoning for CodeLlama
AlgebraicStack Arxiv OpenWebMath GenCode GenLanguage
Baseline 0.189 0.500 0.259 0.029 0.020
Velocitune 0.185 0.463 0.225 0.086 0.040